### v1

# Don’t flatten, tokenize! Unlocking the key to SoftMoE’s efficacy in deep RL

Abstract
The use of deep neural networks in reinforcement learning (RL) often suffers from performance degradation as model size increases. While soft mixtures of experts (SoftMoEs) have recently shown promise in mitigating this issue for online RL, the reasons behind their effectiveness remain largely unknown. In this work we provide an in-depth analysis identifying the key factors driving this performance gain. We discover the surprising result that tokenizing the encoder output, rather than the use of multiple experts, is what is behind the efficacy of SoftMoEs. Indeed, we demonstrate that even with an appropriately scaled single expert, we are able to maintain the performance gains, largely thanks to tokenization.
Figure 1: Comparison of IQM (Agarwal et al., 2021) for Rainbow (Hessel et al., 2018), using the Impala architecture (Espeholt et al., 2018), with penultimate layer scaled, and with SoftMoE with varying numbers of experts, and a single, scaled, expert. IQM scores computed over 200M environment steps across 20 games, with 5 independent runs each. Error bars represent 95% stratified bootstrap confidence intervals. A single scaled expert matches the performance of multiple experts.
1 Introduction
Deep networks have been central to many of the successes of reinforcement learning (RL) since their initial use in DQN (Mnih et al., 2015a), which has served as the blueprint (in particular with regards to network architecture) for most of the followup works. However, recent works have highlighted the need to better understand the network architectures used, and question whether alternate choices can be more effective (Graesser et al., 2022; Sokar et al., 2022, 2023; Obando Ceron et al., 2024). Indeed, Obando Ceron* et al. (2024) and Willi* et al. (2024) demonstrated that the use of Mixtures-of-Experts (MoEs) (Shazeer et al., 2017), and in particular soft mixtures-of-experts (SoftMoEs) (Puigcerver et al., 2024), yields models that are more parameter efficient. Obando Ceron* et al. (2024) hypothesized that the underlying cause of the effectiveness of SoftMoEs was in the induced structured sparsity in the network architecture. While some of their analyses did seem to support this, the fact that performance gains were observed even with a single expert seems to run counter to this hypothesis.
In this work we seek to better understand the true underlying causes for the effectiveness of SoftMoEs in RL settings. Specifically, we perform a careful and thorough investigation into the many parts that make up SoftMoEs aiming to isolate the impact of each. Our investigation led to the surprising finding that the core strength of SoftMoEs lies in the tokenization of the outputs of the convolutional encoder.
The main implications of this finding are two-fold. On the one hand, it suggests that we are still under-utilizing experts when incorporating MoEs in deep RL, given that we are able to obtain comparable performance with a single expert (see Figure 1). On the other hand, this finding has broader implications for deep RL agents trained on pixel-based environments, suggesting that the common practice of flattening the multi-dimensional outputs of the convolutional encoders is ineffective.
Our main contributions can be summarized as follows:
  • We conduct a series of analyses to identify the key factors that allow soft Mixture of Experts (SoftMoEs) to effectively scale RL agents.
  • We demonstrate that tokenization plays a crucial role in the performance improvements.
  • We demonstrate that experts within the SoftMoEs are not specialized to handle specific subsets of tokens and exhibit redundancy. We explore the use of recent techniques for improving network utilization and plasticity to mitigate this redundancy.
The paper is organized as follows. We introduce the necessary background for RL in Section 2 and MoEs in Section 3; we present our experimental setup in Section 4.1, and present our findings from Section 4.2 to Section 6; finally, we discuss related works in Section 7 and provide a concluding discussion in Section 8.
2 Online Deep Reinforcement Learning
In online reinforcement learning, an agent interacts with an environment and adjusts its behaviour based on the reinforcement (in the form of rewards or costs) it receives from the environment. At timestep $t$, the agent inhabits a state $x_{t}\in\mathcal{X}$ (where $\mathcal{X}$ is the set of all possible states) and selects an action $a_{t}\in\mathcal{A}$ (where $\mathcal{A}$ is the set of all possible actions) according to its behaviour policy $\pi:\mathcal{X}\rightarrow\mathfrak{P}(\mathcal{A})$; the environment’s dynamics yield a new state $x_{t+1}\in\mathcal{X}$ and a numerical reward $r_{t}\in\mathbb{R}$. The goal of an RL agent is to find a policy $\pi^{*}$ that maximizes $V^{\pi}(x):=\sum_{t=0}^{\infty}\left[\gamma^{t}r_{t}|x_{0}=x,a_{t}\sim\pi(x_{t})\right]$, where $\gamma\in[0,1)$ is a discount factor. In value-based RL, an agent maintains an estimate of $Q^{\pi}:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}$, defined as $Q^{\pi}(x,a)=\sum_{t=0}^{\infty}\left[\gamma^{t}r_{t}|x_{0}=x,a_{0}=a,a_{t}\sim\pi(x_{t})\right]$; from any state $x\in\mathcal{X}$, a policy $\pi$ can then be improved via $\pi(x)=\arg\max_{a\in\mathcal{A}}Q^{\pi}(x,a)$ (Sutton and Barto, 1998).
In its most basic form, value-based deep reinforcement learning (DRL) extends this approach by using neural networks, parameterized by $\theta$, to estimate $Q$. While the network architectures used may vary across environments, the one originally introduced by Mnih et al. (2015a) for DQN has served as the basis for most. Given that its inputs consist of frames of pixels from Atari games (Bellemare et al., 2013), this network consisted of a series of convolutional layers followed by a series of fully-connected layers; for the remainder of this work we will refer to the convolutional layers as the encoder. The top panel of Figure 2 displays a simplified version of this network, where the encoder produces a 3-dimensional tensor with dimensions $(h,w,d)$; this tensor is flattened to a vector of length $h*w*d$, which is fed through fully-connected layers to produce the estimates for $Q$. One notable variant to the original DQN encoder is Impala’s use of a ResNet encoder (Espeholt et al., 2018), which has been shown to outperform the former. In this work, we mainly focus on the more recent ResNet encoder, but also include some results using the original DQN encoder.
Figure 2: The various architectures considered in this work. Top: baseline architecture; Middle: SoftMoE architecture; Bottom: Top-$k$ architecture.
3 Mixtures of Experts
While there have been a number of architectural variations proposed, in this work we focus on Mixtures of Experts as used by Obando Ceron* et al. (2024). Mixtures of experts (MoEs) were originally proposed as a network architecture for supervised learning tasks using conditional computation, which enabled training very large networks (Shazeer et al., 2017; Fedus et al., 2022). This was achieved by including MoE modules which route a set of input tokens to a set of $n$ sub-networks, or experts; the output of these experts are then combined and fed through the rest of the outer network. The routing/gating mechanism and the expert sub-networks are all trainable. In top-$k$ routing, we consider two methods for pairing tokens and experts. In token choice routing, each token gets an assignation to $k$ experts, each of which has $p$ slots; this may result in certain tokens being dropped entirely by the MoE module and, more critically, certain experts may receive no tokens (Gale et al., 2023). To overcome having load-imbalanced experts, Zhou et al. (2022) propose expert choice routing, where each expert chooses $p$ tokens for its $p$ slots, which results in better expert load-balancing. See the bottom row of Figure 2 for an illustration of this. In most of our analyses, we focus on expert choice routing as it demonstrates higher performance than token choice routing (see Appendix B.1 for full empirical comparison between the two across varying number of experts).
Puigcerver et al. (2024) introduced SoftMoEs, which replace the hard token-expert pairing with a “soft” assignation. Specifically, rather than having each expert slot assigned to a specific token, each slot is assigned a (learned) weighted combination of all the tokens. Thus, each expert receives $p$ weighted averages of all the input tokens. See the middle row of Figure 2 for an illustration of this.
Obando Ceron* et al. (2024) proposed replacing the penultimate dense layer of a standard RL network with an MoE module; the authors demonstrated that, when using SoftMoEs, value-based online RL agents exhibit signs of a “scaling law”, whereby adding more parameters (in the form of more experts) monotonically improves performance. This is in stark contrast with what is observed with the original architectures, where extra parameters result in performance degradation (see Figure 1). The authors argued that the structured sparsity induced by the MoE architecture is what was primarily responsible for their strong performance.
In order to maintain the shape of the input tokens (as originally done with MoE architectures), Obando Ceron* et al. add a projection layer after the “combine” step ($\Phi_{C}$ in Figure 2). Another important detail is the proposed tokenization scheme. MoEs are most commonly used in NLP transformer architectures where tokens are clearly defined (typically as encodings of words), but this tokenization is not something readily available for online RL agents. Thus, Obando Ceron* et al. proposed a number of mechanisms for tokenizing the output of the encoder.
Figure 3: Tokenization schemes considered in this work, specified in the bottom row.
Tokenization
As mentioned in section 2, the output of a standard RL encoder is a 3-dimensional tensor with dimensions $(h,w,d)$; Obando Ceron* et al. (2024) proposed a few possible ways of creating tokens from this output. PerConv tokenization creates $h*w$ tokens of dimension $d$; PerFeat tokenization creates $d$ tokens of dimension $h*w$; see Figure 3 for an illustration. Although necessary for compatibility with transformer MoEs, this tokenization is a departure from the standard practice of flattening the encoder output, as discussed in Section 2.
4 Understanding the impact of the SoftMoE components
While Obando Ceron* et al. (2024) demonstrated the appeal of SoftMoEs for online RL agents, it was not clear whether all of the components of SoftMoEs were important for the performance gains, or whether some were more critical. Further, it was observed that SoftMoEs failed to provide similar gains in actor-critic algorithms such as PPO (Schulman et al., 2017) and SAC (Haarnoja et al., 2018). This section aims to isolate the impact of each of these components.
4.1 Experimental setup
The genesis of our work is to provide a deeper understanding of SoftMoEs in online RL. For this reason, we will mostly focus on SoftMoEs applied to the Rainbow agent (Hessel et al., 2018) with 4 experts and the Impala achitecture (Espeholt et al., 2018), as this was the most performant setting observed by Obando Ceron* et al. (2024). We explore variations of this setting where useful in clarifying the point being argued. We evaluate performance on the same $20$ games of the Arcade Learning Environment (ALE) (Bellemare et al., 2013) benchmark used by Obando Ceron* et al. (2024) for direct comparison, with $5$ independent seeds for each experiment. However, we include the results of our main findings on the full suite of 60 games. We will use the notation “SoftMoE-$n$ to denote a SoftMoE architecture with $n$ experts. All experiments were run on Tesla P100 GPUs using the same Dopamine library (Castro et al., 2018) used by Obando Ceron* et al. (2024); each run with 200 million environment steps took between six and eight days. Following the guidelines suggested by Agarwal et al. (2021), we report human-normalized aggregated interquartile mean (IQM), regular mean, median, and optimality gap, with error bars indicating 95% stratified bootstrap confidence intervals. Performance gains are indicated by increased IQM, median, and mean scores, and reduced optimality gaps. See Appendix A for further details.
4.2 Analysis of SoftMoE Components
At a high level, SoftMoE modules for online RL agents consist of (i) the use of a learnable tensor $\Phi$ used to obtain the dispatch ($\Phi_{D}$) and combine ($\Phi_{C}$) weights; (ii) processing $p$ input slots per expert; (iii) the architectural dimensions (network depth and width); (iv) the use of $n$ experts; and (v) the tokenization of the encoder output. The middle row of Figure 2 provides an illustration of these. In this section we discuss the first four with empirical evaluation demonstrated in Figures 1 and 4, and devote a separate section for tokenization, which turns out to be crucial for performance.
Figure 4: Understanding the impact of SoftMoE components. Using Rainbow with the Impala architecture as our base, we investigate key aspects of SoftMoE-4: (i) combining tokens, (ii) expert specialization, and (iii) adjusting architectural dimensions. SoftMoE does not appear to suffer from the performance degradation observed in the baseline, even when increasing its depth.
Processing combined tokens
To assess the impact of SoftMoE’s weighted combination of tokens via the learnable weights $\Phi$, we compare it with expert choice routing, where experts select sets of (non-combined) tokens. While combined tokens yield improved performance, expert choice routing itself demonstrates substantial gains over the baseline and effectively scales the network (see group [i] in Figure 4). Thus, combined tokens alone do not explain SoftMoE’s efficacy.
Expert specialization
Each expert in a MoE architecture has a limited set of slots: with token/expert selection each slot corresponds to one token, while in SoftMoE each slot corresponds to a particular weighted combination of all tokens. This constraint could encourage each expert to specialize in certain types of tokens in a way that is beneficial for performance and/or scaling. To investigate, we increase the number of slots to be equal to the number of tokens (effectively granting unlimited capacity). Group [ii] in Figure 4 demonstrates that for both SoftMoE and expert choice routing, the performance sees little change when increasing the number of slots, suggesting that expert specialization is not a major factor in SoftMoE’s performance.
Expert width
Obando Ceron* et al. scaled up the penultimate layer of the baseline model with the intent of matching the parameters of the SoftMoE networks, thereby demonstrating that naïvely scaling width degrades performance. While they demonstrated that a performance drop is not observed when scaling down the dimensionality of each expert, they did not investigate what occurs when scaling up the experts dimensionality. This experiment makes sense if comparing the baseline’s penultimate layer dimensionality with that of each expert, rather than with the combined experts dimensionality. Scaling up the width of experts in SoftMoE does not lead to the same performance degradation observed in the scaled baseline model as shown in the first group [iii] of Figure 4.
Network depth
As mentioned in Section 3, SoftMoE models have an additional projection layer within each expert to maintain input token size, resulting in a different number of layers compared to the baseline model. To determine whether this extra layer contributes to SoftMoE’s performance gains, we add a layer of the same size to the baseline model after the penultimate layer. The second group [iii] in Figure 4 demonstrates that this added layer does not improve the baseline model’s performance nor its ability to scale, suggesting that SoftMoE’s performance gains are not attributable to the presence of the extra projection layer.
Multiple experts
Obando Ceron* et al. (2024) demonstrated that multiple experts improves performance. However, the performance gains plateau beyond a certain point, with minimal improvement observed after utilizing four experts (see Figure 1). Consequently, the presence of multiple experts cannot fully account for the observed performance advantages over the baseline model, suggesting that other factors are also contributing to the improved results.
Figure 5: Evaluating impact of tokenizing Rainbow-lite with CNN (top) and Impala (bottom) architectures, using sum/mean over convolutional features. Tokenization with a scaled representation can yield significant performance gains.
5 Don’t flatten, tokenize!
A crucial structural difference remains in how SoftMoE and the baseline model process the encoder’s output: while the baseline model flattens the output of the final convolutional layer, SoftMoE divides it into distinct tokens (see Section 3). We conduct a series of experiments to properly assess the impact of this choice, going from the standard baseline architecture to a single-expert SoftMoE.
Tokenized baseline
Our first experiment in this series simply adds tokenization to the baseline architecture. Specifically, given the 3-dimensional encoder output with shape $[h,w,d]$, we replace the standard flattening (to a vector of dimension $h*w*d$) with a reshape to a matrix of shape $[h*w,d]$; then, prior to the final linear layer, we either sum over the first dimension or take their average. Since this operation is not immediately applicable to Rainbow (due to Noisy networks), we explored this experiment with Rainbow-lite, a version of Rainbow in Dopamine which only includes C51 (Bellemare et al., 2017), multi-step returns, and prioritized experience replay. Figure 5 demonstrates that this simple change can yield statistically significant improvements for Rainbow-lite, providing strong evidence that tokenization plays a major role in the efficacy of SoftMoE.
Single expert processing all tokens sparsely
To assess the influence of tokenization within the MoE framework, while ablating all other four components studied in Section 4.2, we explore using a single expert to process all input tokens. Each token is sparsely processed in one of the single expert slots. This corresponds to expert choice-1. In line with the “expert width” investigation from the last section, we study expert choice-1 with the scaled expert dimensionality to match the scaled baseline one. As Figure 6 shows, even with a single expert the performance is comparable to that of SoftMoE with multiple experts. This suggests that tokenization is one of the major factors in the performance gains achieved by SoftMoE.
Figure 6: Combined tokenization is a key driver of SoftMoE’s efficacy. SoftMoE with a single scaled expert has comparable performance with using multiple experts. Even non-combined tokenization (e.g. ExpertChoice) with a single expert achieves similar performance.
Combined tokenization is a key driver
We further investigate the impact of combined tokenization in the single expert setting. To this end, we study SoftMoE-1, with scaled expert dimensionality as before. As Figure 6 shows, using combined tokens with a single expert brings the performance even closer to the multiple expert setting. This further confirms that (combined) tokenization is the key driver of SoftMoE’s efficacy.
5.1 Additional analyses
The main takeaway from the previous section is that SoftMoE with a single (scaled) expert is as performant as the variant with multiple unscaled experts. To better understand this finding, in this section we perform an extra set of experiments with this variant of SoftMoE.
Figure 7: Comparison of different tokenization schemes for SoftMoE-1 with the standard and scaled penultimate layer. PerConv achieves the highest performance, suggesting that maintaining the spatial structure of the features plays a role on the observed benefits.
Figure 8: Evaluating the impact of fewer slots with a single scaled (by 4x) expert for both SoftMoE and ExpertChoice. Even with drastically fewer slots, SoftMoE-1 remains competitive.
Are all types of tokenization equally effective?
We compare the different tokenization types presented in Section 3, in addition to two new types of tokenization, which serve as useful references: PerPatch tokenization splits the 3-dimensional $[h,w,d]$ tensor into $\rho$ patches, over which average pooling is performed resulting in a new tensor of dimensions $[h/\rho,w/\rho,d]$; Shuffled is similar to PerConv, except it first applies a fixed permutation on the encoder output. See Figure 3 for an illustration of these. The comparison of the different tokenization types is summarized in Figure 7. Consistent with the findings of Obando Ceron* et al. (2024), PerConv performs best, and in particular outperforms PerFeat. It is interesting to observe that PerPatch is close in performance to PerConv, while Shuffled is the worst performing. The degradation from PerPatch is likely due to the average pooling and reduced dimensionality; the degradation from shuffling is clearly indicating that it is important to maintain the spatial structure of the encoding output. Combining these results with our previous findings suggests that using tokenization maintains the spatial structure learned by the convolutional encoder, and is crucial to maximizing performance.
Computational efficiency by combined tokens
Puigcerver et al. (2024) highlighted that the time complexity of SoftMoE depends directly on the number of slots. As can be seen in Figure 8, using SoftMoE-1 with only 10% of the default number of slots, and even with a single slot, achieves a comparable performance to using the regular number of slots. On the other hand, the reduced performance of ExpertChoice-1 further confirms the gains added by combining tokens in SoftMoE.
Figure 9: (Left) Evaluating the performance on all the 60 games of Atari 2600. Consistent with our results SoftMoE-1 matches the performance of multiple experts over the full suite. (Right) CNN Encoder. The performance gain of SoftMoE-1 is present when training with the standard CNN encoder, revealing the importance of tokenization across architectures.
More games and different encoders
Our results so far were run on the 20 games used by Obando Ceron* et al. (2024) and with the Impala ResNet encoder (Espeholt et al., 2018). To ensure our findings are not specific to these choices, we extend our evaluation to all 60 games in the ALE suite, and with the standard CNN encoder of Mnih et al. (2015b). Our findings, as illustrated in Figure 9, demonstrate consistent results across all games and various encoder architectures.
Figure 10: Evaluating on DQN (top) and DER (bottom). When combined with DQN, SoftMoE with a single scaled expert does not provide gains over the baseline. In contrast, when combined with DER, it achieves performance comparable to SoftMoE-4.
More agents
We focused on Rainbow for the majority of our investigation due to the positive results demonstrated by Obando Ceron* et al. (2024), but it is important to also consider other agents. To this point, we evaluate the impact of a single scaled expert on DQN (Mnih et al., 2015b). Note that we replace the original Huber loss used in DQN with mean-squared error loss, as it has been shown to be more performant (Ceron and Castro, 2021), and coincides with what was used by Obando Ceron* et al.. and DER (Van Hasselt et al., 2019). Following the training regime used by Obando Ceron* et al. (2024), we train DER for 50M environment frames. In the DQN setting, SoftMoE-1 does not appear to offer significant improvements over the baseline, potentially due to the limited gains observed with SoftMoE-4 (Figure 10, top). Further investigation is needed to understand how the inherent instability of DQN might limit the potential benefits of SoftMoE. On the other hand, consistent with our finding in Rainbow, in DER, where SoftMoE-4 achieves a superior performance over the baseline, SoftMoE-1 is able to offer the same gains (Figure 10, bottom).
6 Calling all experts!
Our findings in Sections 4 and 5 reveal that the main benefit of using SoftMoE architectures is in the tokenization and their weighted combination, rather than in the use of multiple experts. The fact that we see no evidence of expert specialization suggests that most experts are redundant (see Figure 4). Rather than viewing this as a negative result, it’s an opportunity to develop new techniques to improve expert utilization and maximize the use of MoE architectures in RL.
As a first step in this direction, we adopt existing techniques designed to improve network utilization: (1) periodic reset of parameters (Nikishin et al., 2022) and (2) using Shrink and Perturb (S&P) (Ash and Adams, 2020), as used by D’Oro et al. (2022) and Schwarzer* et al. (2023). We perform these experiments on Rainbow and DER with SoftMoE-4. We explore applying resets and S&P to all experts and to a subset of experts, where the experts to reset are those that have the highest fraction of dormant neurons (Sokar et al., 2023). For Rainbow, we reset every 20M environment steps, whereas for DER we reset every 200k environment steps.
Interestingly, the effectiveness of these techniques varied across different reinforcement learning agents. While weight reset and S&P did not improve expert utilization in the Rainbow agent, they led to enhanced performance in the sample-efficient DER agent (Figure 11). Within the DER agent, resetting a subset of experts proved more beneficial than resetting all, while applying S&P to all experts yielded the most significant performance gains. These findings highlight the need for further research into methods that can optimize expert utilization and improve performance in reinforcement learning with SoftMoE.
Figure 11: Evaluating the use of resets (top) and S&P (bottom) for improving expert utilization on Rainbow (left column) and DER (right column). All experiments run with SoftMoE-4. Resets and S&P seem to benefit DER, but degrade performance for Rainbow.
7 Related Works
Mixture of Experts
Mixture of Experts (MoEs) have demonstrated remarkable success in scaling language and computer vision models to trillions of parameters (Lepikhin et al., 2020; Fedus et al., 2022; Wang et al., 2020; Yang et al., 2019; Abbas and Andreopoulos, 2020; Pavlitskaya et al., 2020). Chung et al. (2024) examined the implicit biases arising from token combinations in SoftMoE and defined the notion of expert specialization in computer vision. Besides the success of MoEs in single task training, similar gains have been observed in transfer and multitask learning (Puigcerver et al., 2020; Chen et al., 2023; Ye and Xu, 2023; Hazimeh et al., 2021). In addition to the works of Obando Ceron* et al. and Willi* et al. already discussed, Farebrother et al. (2024) proposed replacing regression loss with classification loss and showed its benefits on MoEs.
Network Scalability in RL
Scaling reinforcement learning (RL) networks presents significant challenges, largely due to issues with training instabilities (Hessel et al., 2018). To address these difficulties, several techniques have been proposed, with a focus on algorithmic improvements or hyperparameter optimization (Farebrother et al., 2022, 2024; Taiga et al., 2022; Schwarzer* et al., 2023). Recently, a few studies have begun to focus on architectural innovations as a means of addressing these challenges. Kumar et al. (2023a) showed that incorporating a learned spatial embedding with the output of a ResNet encoder, prior to flattening it, improves the scalability in offline RL. Obando Ceron et al. (2024) demonstrated that gradual magnitude pruning enables the scaling of value-based agents, resulting in substantial performance gains.
Network Plasticity in RL
It has been observed that networks trained on non-stationary distributions tend to lose their expressivity and ability to adapt to new data over time, a phenomenon referred to as the loss of plasticity (Kumar et al., 2021; Lyle et al., 2021). While several studies have attempted to explain this phenomenon (Lyle et al., 2023; Ma et al., 2023; Lyle et al., 2024), the precise underlying causes remain uncertain. In a parallel line of research, numerous methods have been proposed to mitigate plasticity loss and improve network trainability. These approaches involve using small batch size for gradient updates (Ceron et al., 2023), periodically resetting some of the network’s parameters to random values (Sokar et al., 2023; Nikishin et al., 2022; Igl et al., 2020; Dohare et al., 2024), slightly modifying the existing weights by shrinking and perturbing them (D’Oro et al., 2022; Schwarzer* et al., 2023), and restricting the large deviation from initial weights (Kumar et al., 2023b; Lewandowski et al., 2024). Recently, Lee et al. have also demonstrated that combining reinitialization with a teacher-student framework can enhance plasticity. Another technique involves preserving useful neurons by perturbing gradients to ensure smaller updates for important neurons while allowing larger updates for less important ones (Elsayed and Mahmood, ). In this work, we explore some of these methods to improve plasticity and utilization of experts.
8 Conclusions
We began our investigation by evaluating the components that make up a SoftMoE architecture, so as to better understand the strong performance gains reported by Obando Ceron* et al. (2024) when applied to value-based deep RL agents. One of the principal, and rather striking, findings is that tokenizing the encoder outputs is one of the primary drivers of the performance improvements. Since tokenization is mostly necessary for pixel-based environments (where encoders typically consist of convolutional layers), it can explain the absence of performance gains observed by Obando Ceron* et al. (2024) and Willi* et al. (2024) when using MoE architectures in actor-critic algorithms, since these were evaluated in non-pixel-based environments.
More generally, our findings suggests that the common practice of flattening the outputs of deep RL encoders is sub-optimal, as it is likely not maintaining the spatial structure output by the convolutional encoders. Indeed, the results with PerPatch and Shuffled tokenization in Figure 7 demonstrate that maintaining this spatial structure is important for performance. This is further confirmed by our results in Figure 5, where we can observe that not flattening the convolutional outputs can lead to significant improvements. This idea merits further investigation, as the results vary depending on the architecture used and whether features are summed or averaged over.
It is somewhat surprising that our results do not seem to carry over to DQN. However, the fact that we do see similar results with Rainbow-lite and DER (which is based on Rainbow) suggests that the use of categorical losses may have a complementary role to play. This is a point argued by Farebrother et al. (2024) and merits further inquiry.
Although the use of mixtures-of-experts architectures for RL show great promise, our results suggest that most of the experts are not being used to their full potential. The findings in Figure 11 present one viable mechanism for increasing their use, but further research is needed to continue increasing their efficacy for maximally performant RL agents.
Acknowledgements
The authors would like to thank Gheorghe Comanici, Jesse Farebrother, Joan Puigcerver, Doina Precup, and the rest of the Google DeepMind team for valuable feedback on this work. We would also like to thank the Python community (Van Rossum and Drake Jr, 1995; Oliphant, 2007) for developing tools that enabled this work, including NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007) and JAX (Bradbury et al., 2018).
Appendix A Experimental Details
Tasks
We evaluate Rainbow on 20 games from the Arcade Learning Environment (Bellemare et al., 2013). We use the same set of games evaluated by Obando Ceron* et al. (2024) for direct comparison. This includes: Asterix, SpaceInvaders, Breakout, Pong, Qbert, DemonAttack, Seaquest, WizardOfWor, RoadRunner, BeamRider, Frostbite, CrazyClimber, Assault, Krull, Boxing, Jamesbond, Kangaroo, UpNDown, Gopher, and Hero. In addition, we run some of our main results on all suite of $60$ games.
On Atari100k benchmark, we evaluate on the full $26$ games. These are: Alien, Amidar, Assault, Asterix, BankHeist, BattleZone, Boxing, Breakout, ChopperCommand, CrazyClimber, DemonAttack, Freeway, Frostbite, Gopher, Hero, Jamesbond, Kangaroo, Krull, KungFuMaster, MsPacman, Pong, PrivateEye, Qbert, RoadRunner, Seaquest, UpNDown.
Hyper-parameters
We use the default hyper-parameter values for DQN (Mnih et al., 2015a), Rainbow (Hessel et al., 2018), and DER (Van Hasselt et al., 2019) agents. We share the details of these values in Table 1.
Table 1: Default hyper-parameters setting for DQN, Rainbow and DER agents.
Atari
Hyper-parameter DQN Rainbow DER
Adam’s ($\epsilon$) 1.5e-4 1.5e-4 0.00015
Adam’s learning rate 6.25e-5 6.25e-5 0.0001
Batch Size 32 32 32
Conv. Activation Function ReLU ReLU ReLU
Convolutional Width 1 1
Dense Activation Function ReLU ReLU ReLU
Dense Width 512 512 512
Normalization None None None
Discount Factor 0.99 0.99 0.99
Exploration $\epsilon$ 0.01 0.01 0.01
Exploration $\epsilon$ decay 250000 250000 2000
Minimum Replay History 20000 20000 1600
Number of Atoms 0 51 51
Number of Convolutional Layers 3 3 3
Number of Dense Layers 2 2 2
Replay Capacity 1000000 1000000 1000000
Reward Clipping True True True
Update Horizon 1 3 10
Update Period 4 4 1
Weight Decay 0 0 0
Sticky Actions True True False
Appendix B Additional Experiments
B.1 Comparison between token choice routing and expert choice routing
We focus in our analysis on expert choice routing, where experts select tokens, as opposed to token choice routing, where tokens select experts. We compare these two routing strategies across different numbers of experts using a Rainbow agent with a ResNet architecture.
Figure 12 demonstrates that expert choice routing consistently outperforms token choice routing across all numbers of experts. Moreover, this performance gap widens as the number of experts increases, highlighting the advantages of expert choice routing in addressing the limitations of token choice routing.
Figure 12: A comparison between token choice routing and expert choice routing on Rainbow with ResNet encoder. Expert choice routing consistently outperforms token choice routing across different number of experts. We report interquantile mean performance with error bars indicating 95% confidence intervals, over 20 games with 5 seeds each.
B.2 Don’t flatten tokenize
In this section, we provide further analysis on the impact of tokenization and their weight combination on models with larger number of experts. To this end, we investigate SoftMoE-8 and compare it against the scaled SoftMoE-1 and ExpertChoice-1 (non-combined tokenization). Figure 13 shows that single scaled expert can reach a performance close to 8 experts. In addition, consistent with our results, combined tokenization adds benefits over the sparse one. Figure 14 shows a comparison between the learning curves of single scaled expert and multiple ones.
Figure 13: Combined tokenization plays a crucial role in performance. SoftMoE-1 with scaled hidden layer achieves a comparable performance to SoftMoE-8.
Figure 14: Comparison between the learning curves of the scaled SoftMoE-1 and multiple experts. We report interquantile mean performance with error bars indicating 95% confidence intervals, over 20 games with 5 seeds each.
B.3 More agents
Same as our finding for DQN using SoftMoE-4, we observe that single expert has a closer performance to the scaled baseline than SoftMoE-8, as shown in Figure 15. Further investigation is needed to fully understand the behavior of SoftMoE on DQN.
Figure 15: Evaluating more agents. When combined with DQN, SoftMoE with a single scaled expert does not provide large gains over the baseline.
B.4 Expert Utilization
In this section we provide more analysis to understand the utilization of experts in multi expert setting and the effect of existing techniques on improving plasticity.
Some experts are redundant
In Section 4.2, we showed that expert are not specialized in learning subset of tokens. We expand this analysis by assessing the redundancy of experts. To this end, we prune two experts of SoftMoE-4 during training. We studied two pruning schemes: pruning the two experts once and gradually prune throughout training. As shown in Figure 16, pruning these experts does not change the performance of the agent.
Figure 16: Pruning subset of experts in SoftMoE-4 does not lead to performance drop, suggesting that some of the experts are redundant. We report interquantile mean performance with error bars indicating 95% confidence intervals, over 20 games with 5 seeds each.
Reset experts and router weights
In investigating current techniques to improve expert utilization, we explore reset either experts weights only or both experts and router weights. As illustrated in Figure 17, both methods do not help in improving utilization, with resetting the router weights leads to worse performance.
Figure 17: Comparison between resetting experts weights only versus experts and router weights in Rainbow (top) and DER (bottom).